OpenStudy (anonymous):

help me find the greatest common factor gcf and the least common factor (lcm) i'm just in grade two

OpenStudy (anonymous):

give me an example

OpenStudy (anonymous):

in general: if the number ends with 0 or 5, you may divide by 5; if ends by 0,2,4,6,8, you may divide by 2; if the number like this abc, for example 210, a+b+c can be divided by 3 (2+1+0=3 can be divided by 3), then the number abc can be divided by 3. Then you put all the divided number in order,choose the same number

OpenStudy (anonymous):

LCM is the smallest number of all the numbers, and GCM the largest number means all the same numbers multiply together

OpenStudy (anonymous):

for example: 48 and 32, 48 and 32 can be both divided by 2, so 2 is the LCM; 48 and 32 can be divided by 16, so 16 is the GCM; 32=16*2,48=16*3